# Xu Shen  and  Yuqiang Zheng Morningside Center of Mathematics Academy of Mathematics and Systems Science Chinese Academy of Sciences

### v2

@@ 5.4. Extensions of $\mathcal{G}_{0}^{\mathrm{spl}}$-zips and Hasse invariants
$\mathfrak{X}_{\sigma}^{\mathrm{spl},\circ}\simeq(\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}})^{\wedge}_{\cup_{\tau\in\Sigma_{Z}^{+},\overline{\tau}\subset\overline{\sigma}}Z_{[\tau]}}.$
@@
$\mathbf{W}\rightarrow\mathsf{M}_{K,\Sigma,0}^{\mathrm{spl},\mathrm{tor}},\quad\mathbf{W}\rightarrow\Xi^{\mathrm{spl}}(\sigma).$
By these morphisms, the two stratifications of $W$ induced respectively by those of $\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}$ and $\Xi^{\mathrm{spl}}(\sigma)$ coincide. Let
@@
be the open stratum, which is the preimages of $\mathsf{M}_{K}^{\mathrm{spl}}$ and $\Xi^{\mathrm{spl}}$ under the above morphisms. Now consider the geometric special fiber $\mathsf{M}_{K,\Sigma,0}^{\mathrm{spl},\mathrm{tor}}$ and we denote by the same notations $Z^{\mathrm{spl}},C^{\mathrm{spl}},\Xi^{\mathrm{spl}},\mathbf{W}$ etc the corresponding objects base-changed to $k$. By [36] Definition 2.2.1, a locally closed subset
@@
the preimages of $Y$ and $Y_{Z}^{\sharp}$ in $W^{0}$ coincide. Here $\mathbf{W}^{0}\rightarrow Z^{\mathrm{spl}}$ is the composition
@@ 6.1. Hecke actions on coherent cohomology
the restricted tensor product of the spherical Hecke algebras $\mathcal{H}_{v}=\mathbb{Z}_{p}[K_{v}\backslash G(\mathbb{Q}_{p})/K_{v}]$ outside $S$. We assume that $G$ is ramified over $\mathbb{Q}_{p}$, i.e. $p\in S$, since otherwise all the following discussions are covered by [18].
@@ 6.2. Factorizations to $H^{0}$
    There exists an arithmetic progression $A$ such that $\eta+a\eta_{\omega}\in F(i,n,\eta)$ for all $a\in A\cap\mathbb{Z}_{\geq 1}$.
@@
together with some integers $a_{j}$ and global sections $f_{j}\in H^{0}(Z_{j},\omega_{\operatorname{Hdg}}^{a_{j}})$, such that each $f_{j}$ is a lifting of a length Hasse invariant, and for each $10\leq j\leq r-$ 1, we have $Z_{j+1}=V(f_{j})$. More precisely, on the reduced locus
@@
By Proposition 4.4, there exists a large integer $N_{d-j}$ and a length $d-j$ Hasse invariant
@@ 6.3. Galois representations
where $R({}^{L}G_{v})$ is the algebra obtained by restricting characters representations of ${}^{L}G_{v}$ to semisimple ${}^{L}G_{v}^{\circ}(\overline{\mathbb{Q}}_{p})$-conjugacy classes in ${}^{L}G_{v}\rtimes\operatorname{Frob}_{v}$. If $\pi_{v}$ is an unramified irreducible smooth representation of $G(\mathbb{Q}_{v})$, we get the corresponding semisimple ${}^{L}G_{v}^{\circ}(\overline{\mathbb{Q}}_{p})$-conjugacy class $\mathrm{Sat}(\pi_{v})$, the Satake parameter of $\pi_{v}$.
@@ A.1. Integral models of PEL-type Shimura varieties
We keep the notations of section 2. Let $G$ be the group defined by the rational PEL datum and $\mathscr{L}$ a multichain of $\operatorname{\mathcal{O}}_{B}$-modules. By [47], there are three integral models $\mathscr{A}^{\operatorname{naive}},\mathscr{A}$ and $\mathscr{A}^{\mathrm{spl}}$ of the PEL moduli space over $E$ (with respect to the multichain $\mathscr{L}$). The models $\mathscr{A}^{\operatorname{naive}},\mathscr{A}$ are defined over $\operatorname{\mathcal{O}}_{E}$ and $\mathscr{A}^{\mathrm{spl}}$ is defined over $\operatorname{\mathcal{O}}_{F}$. We first recall the definition of PEL datum with parahoric level structure following [53], see also the appendix of [57] or [21] section 2.
@@ A.1.1. Parahoric data at $p$
    A chain of $\operatorname{\mathcal{O}}_{B}$-lattices in $V$ is a set of totally ordered $\operatorname{\mathcal{O}}_{B}$-lattices $\mathscr{L}$ such that for every element $x\in B^{\times}$ which normalizes $\operatorname{\mathcal{O}}_{B}$, one have
@@
for the group scheme over $\mathbb{Z}_{p}$ defined by the stabilizer of multichain $\mathscr{L}$ as in [53]*§6. By [19]*Corollary 4.8, for a reductive group of the form $G=\operatorname{Res}_{F|\operatorname{\mathbb{Q}}_{p}}G^{\prime}$, every parahoric group scheme of $G$ is of the form $\operatorname{Res}_{\operatorname{\mathcal{O}}_{F}|\operatorname{\mathbb{Z}}_{p}}\operatorname{\mathcal{G}}^{\prime}$ for a unique parahoric group scheme $\operatorname{\mathcal{G}}^{\prime}$ of $G^{\prime}$. If we assume that $G^{\prime}$ unramified, the stabilizer group scheme $\operatorname{\mathcal{G}}=\operatorname{\mathcal{G}}_{\mathscr{L}}$ associated to $\mathscr{L}$ is always connected. So $\operatorname{\mathcal{G}}$ is a parahoric group scheme of $G$ and (since the prime to $p$ level $K^{p}$ is fixed) we write $K=\mathcal{G}(\mathbb{Z}_{p})$.
@@ A.1.2. Integral models
[…]
@@
  • $\overline{\eta}$ is a $\pi_{1}(S,\overline{s})$-invariant $K^{p}$-orbit of isomorphism $\eta:V\otimes\mathbb{A}_{f}^{p}\to T^{p}(A_{\overline{s}})$ which preserve the pairings up to a scalar in $(\mathbb{A}_{f}^{p})^{\times}$. Here $T^{p}(A_{\overline{s}})$ is the prime to $p$ Tate module of $A_{\overline{s}}$ and for simplicity we assume that $S$ is connected.
@@
Suppose that $S$ is a scheme over $\operatorname{\mathcal{O}}_{F}$. A $\mathscr{L}$-set of polarized $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{S}$-modules is a triple (¯,¯$(\underline{\operatorname{\mathcal{H}}},\underline{\mathscr{F}},\underline{\iota})$, where:
@@
    ¯$\underline{\operatorname{\mathcal{H}}}:\Lambda\mapsto\operatorname{\mathcal{H}}_{\Lambda}$ and ¯$\underline{\mathscr{F}}:\Lambda\mapsto\mathscr{F}_{\Lambda}$ are functors from the category $\mathscr{L}$ to the category of $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{S}$-modules.
@@
    For each $\Lambda\in\mathscr{L}$, both $\mathscr{F}_{\Lambda}$ and $\operatorname{\mathcal{H}}_{\Lambda}/\mathscr{F}_{\Lambda}$ are finite locally free $\operatorname{\mathcal{O}}_{S}$-modules, and that $\operatorname{\mathcal{H}}_{\Lambda}/\mathscr{F}_{\Lambda}$ satisfies the determinant condition.
@@
Suppose that $S$ is a scheme over $\operatorname{\mathcal{O}}_{F}$, and that (¯,¯$(\underline{\mathcal{H}},\underline{\mathscr{F}},\underline{\iota})$ is a $\mathscr{L}$-set of polarized $\operatorname{\mathcal{O}}_{B}\otimes\mathcal{O}_{S}$-modules. A splitting structure for (¯,¯$(\underline{\mathcal{H}},\underline{\mathscr{F}},\underline{\iota})$ is a collection
¯$\underline{\operatorname{\mathscr{F}}_{\bullet}}=(\mathscr{F}_{i,j}^{l},\iota_{i,j}^{l})_{1\leq i\leq r,1\leq j\leq f_{i},0\leq l\leq e_{i}},$
where each $\mathscr{F}_{i,j}^{l}:\Lambda\to\mathscr{F}_{\Lambda,i,j}^{l}$ is a functor from the category $\mathscr{L}$ to the category of $\operatorname{\mathcal{O}}_{B}\otimes\mathcal{O}_{S}$-modules, and each $\iota_{i,j}^{l}:\mathscr{F}_{\Lambda,i,j}^{l}\to\operatorname{\mathcal{H}}_{\Lambda,i,j}$ is an injective morphism satisfies the following conditions (identify $\mathscr{F}_{i,j}^{l}$ with its image under $\iota$):
@@
    For each $\Lambda\in\mathscr{L}$, we require that both $\mathscr{F}_{\Lambda,i,j}^{l}$ and $\operatorname{\mathcal{H}}_{\Lambda,i,j}/\mathscr{F}_{\Lambda,i,j}^{l}$ to be finite locally free $\mathcal{O}_{S}$-modules.
@@
    For each $\Lambda\in\mathscr{L}$ and tuples $(i,j,l)$, let $(\mathscr{F}_{\Lambda,i,j}^{l})^{\perp}$ denote the orthogonal complement of $\mathscr{F}_{\Lambda,i,j}^{l}$ in $\operatorname{\mathcal{H}}_{\Lambda^{\vee},i,j}^{l}$ with respect to the perfect pairing $\operatorname{\mathcal{H}}_{\Lambda,i,j}\times\operatorname{\mathcal{H}}_{\Lambda^{\vee},i,j}\to\operatorname{\mathcal{O}}_{S}$. Then
@@
Let $\mathscr{A}^{\mathrm{spl}}_{\mathscr{L}}$ be the moduli scheme over $\operatorname{\mathcal{O}}_{F}$ classifying the objects $(A_{\mathscr{L}},\overline{\lambda},\overline{\eta},\underline{\mathscr{F}_{\bullet}})$ for each scheme $S/\operatorname{\mathcal{O}}_{F}$, where $(A_{\mathscr{L}},\overline{\lambda},\overline{\eta})$ is an object in $\mathscr{A}^{\operatorname{naive}}_{\mathscr{L}}(S)$ and $\underline{\mathscr{F}_{\bullet}}$ is a splitting structure for the $\mathscr{L}$-set of polarized $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{S}$-modules associated with $(A_{\mathscr{L}},\overline{\lambda},\overline{\eta})$. This is the splitting model associated to the PEL datum. By the notation of [35], we have
@@ A.2. Local models
The local structure of integral models of Shimura varieties are controlled by the associated local models. There is a naive local model $\mathbb{M}^{\operatorname{naive}}$ associated to the moduli scheme $\mathscr{A}^{\operatorname{naive}}$. Recall that $\operatorname{\mathbb{M}}^{\operatorname{naive}}:=\operatorname{\mathbb{M}}^{\operatorname{naive}}(\mathcal{G},\mu)$ is the moduli scheme over $\operatorname{\mathcal{O}}_{E}$ given by the following definition.
@@
    A functor from the category $\operatorname{\mathscr{L}}$ to the category of $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{S}$-modules on $S$:
@@
We require that the following conditions to be satisfied:
@@
    The morphism $\varphi_{\Lambda}$ are surjective.
@@
as moduli scheme over $\operatorname{\mathcal{O}}_{F}$ classifying the splitting structures over $\mathbb{M}^{\operatorname{naive}}\otimes_{\operatorname{\mathcal{O}}_{E}}\operatorname{\mathcal{O}}_{F}$. The splitting model $\mathbb{M}^{\mathrm{spl}}$ admits good properties.
@@
Such multichain determine a parahoric subgroup of $G^{\prime}_{\tau}$. This gives a local model (over $\operatorname{\mathcal{O}}_{F}$)
@@
By our assumption, $G^{\prime}$ is unramified over $F$, so $\operatorname{\mathbb{M}}^{\mathrm{loc}}$ is the same as naive local model. This means that for each $\operatorname{\mathcal{O}}_{F}$-scheme $S$, $\operatorname{\mathbb{M}}^{\mathrm{loc}}(S)$ classifies the set of multichain $\{\operatorname{\mathscr{F}}_{\Lambda,\tau}\}_{\Lambda\in\operatorname{\mathscr{L}}}\subset\operatorname{\mathscr{L}}_{\tau}$ which is compatible with transition maps, and each $\operatorname{\mathscr{F}}_{\Lambda,\tau}$ is Zariski locally on $S$ an $\operatorname{\mathcal{O}}_{S}$-direct summand of $\Lambda_{\tau}$ of rank $d_{\tau}$.
@@
Such a trivialization exists Zariski locally on $S$ by [47]*Propositions 5.2, 9.2 (this reference only cover the case of type (C) and (AL); the case of type (AU) can be proved by the same argument). The map $\pi_{1}$ is the natural forgetful morphism, and $\pi_{2}$ sends $(\operatorname{\mathscr{F}}_{\Lambda,j}^{l},{\varphi_{\Lambda,j}^{l}})$ to $\varphi_{\Lambda,j}^{l}(\operatorname{\mathscr{F}}_{\Lambda,j}^{l}/\operatorname{\mathscr{F}}_{\Lambda,j}^{l-1})\in\operatorname{\mathbb{M}}^{\mathrm{loc}}_{\sigma_{j}^{l}}$.
Let $\operatorname{\mathcal{G}}_{j}^{l}$ be the subgroup of $\prod_{\Lambda\in\operatorname{\mathscr{L}}}\operatorname{Aut}(\Lambda_{j}^{l})$ compatible with the transition maps of $\operatorname{\mathscr{L}}$, and $\operatorname{\mathcal{G}}^{\mathrm{spl}}:=\prod_{j,l}\operatorname{\mathcal{G}}_{j}^{l}$. Then the action
@@
    ${{\widetilde{\mathscr{A}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{q}$${{\mathscr{A}}}$${{\operatorname{\mathbb{M}}^{\mathrm{loc}}}}$
@@
We only need to show $\operatorname{\mathbb{M}}$ coincides with $\operatorname{\mathbb{M}}^{\rm loc}$, as the other statements are contained in [47]. The image $\operatorname{\mathbb{M}}^{\rm loc}$ is flat by the flatness of $\operatorname{\mathbb{M}}^{\mathrm{spl}}$. As the generic fiber of $\operatorname{\mathbb{M}}$ agrees with $\operatorname{\mathbb{M}}^{\operatorname{naive}}$, $\operatorname{\mathbb{M}}$ is the flat closure of the generic fiber of $\operatorname{\mathbb{M}}^{\operatorname{naive}}$, which agrees with $\operatorname{\mathbb{M}}^{\rm\mathrm{loc}}$ by definition. ∎
@@
For each $i,j,l$, let $\operatorname{\mathcal{G}}_{i,j}^{l}$ be the group scheme over $\operatorname{\mathcal{O}}_{F}$ defined by the automorphism of multichain $(\Lambda_{i,j}^{l})_{\Lambda\in\operatorname{\mathscr{L}}}$. Then it admits a decomposition as $\operatorname{\mathcal{G}}^{\mathrm{spl}}=\prod_{i,j,l}\mathcal{G}^{l}_{i,j}$. Now we define
@@
where $\pi^{\prime\prime}$ is a $\operatorname{\mathcal{G}}_{\operatorname{\mathcal{O}}_{F}}\times\prod_{i,j,l\geq 2}\operatorname{\mathcal{G}}_{i,j}^{l}$-torsor, and we can check that the same action of $\pi^{\prime\prime}$ makes $q^{\prime\prime\prime}$ a $\operatorname{\mathcal{G}}_{\operatorname{\mathcal{O}}_{F}}\times\prod_{i,j,l\geq 2}\operatorname{\mathcal{G}}_{i,j}^{l}$-equivariant smooth morphism. Note that we always have $\Lambda_{i}^{1}\subset\Lambda$, so the restriction of $\operatorname{\mathcal{G}}$-action on $\Lambda$ induces a natural morphism

### v3

@@ 5.4. Extensions of $\mathcal{G}_{0}^{\mathrm{spl}}$-zips and Hasse invariants
$\mathfrak{X}_{\sigma}^{\mathrm{spl},\circ}\simeq(\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}})^{\wedge}_{\quad\cup_{\tau\in\Sigma_{Z}^{+},\overline{\tau}\subset\overline{\sigma}}Z_{[\tau]}}.$
@@
$\mathbf{W}\rightarrow\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}},\quad\mathbf{W}\rightarrow\Xi^{\mathrm{spl}}(\sigma).$
By these morphisms, the two stratifications of $\mathbf{W}$ induced respectively by those of $\mathsf{M}_{K,\Sigma}^{\mathrm{spl},\mathrm{tor}}$ and $\Xi^{\mathrm{spl}}(\sigma)$ coincide. Let
@@
be the open stratum, which is the preimage of $\mathsf{M}_{K}^{\mathrm{spl}}$ and $\Xi^{\mathrm{spl}}$ under the above morphisms. Now consider the geometric special fiber $\mathsf{M}_{K,\Sigma,0}^{\mathrm{spl},\mathrm{tor}}$ and we denote by the same notations $Z^{\mathrm{spl}},C^{\mathrm{spl}},\Xi^{\mathrm{spl}},\mathbf{W}$, etc. the corresponding objects base-changed to $k$. By [38] Definition 2.2.1, a locally closed subset
@@
the preimages of $Y$ and $Y_{Z}^{\sharp}$ in $\mathbf{W}^{0}$ coincide. Here $\mathbf{W}^{0}\rightarrow Z^{\mathrm{spl}}$ is the composition
@@ 6.1. Hecke actions on coherent cohomology
[…]
@@ 6.2. Factorizations to $H^{0}$
    There exists an arithmetic progression AA such that $\eta+a\eta_{\omega}\in F(i,n,\eta)$ for all $a\in A\cap\mathbb{Z}_{\geq 1}$.
@@
together with some integers $a_{j}$ and global sections $f_{j}\in H^{0}(Z_{j},\omega_{\operatorname{Hdg}}^{a_{j}})$, such that each $f_{j}$ is a lifting of a length Hasse invariant, and for each $0\leq j\leq r-1$, we have $Z_{j+1}=V(f_{j})$. More precisely, on the reduced locus
@@
By Proposition 4.4, there exists a large integer $N_{d-j}$ and a length djd-j Hasse invariant
@@ 6.3. Galois representations
where $R({}^{L}G_{v})$ is the algebra obtained by restricting character representations of $^{L}G_{v}$ to semisimple $^{L}G_{v}^{\circ}(\overline{\mathbb{Q}}_{p})-$conjugacy classes in $^{L}G_{v}\rtimes\operatorname{Frob}_{v}$. If $\pi_{v}$ is an unramified irreducible smooth representation of $G(\mathbb{Q}_{v})$, we get the corresponding semisimple $^{L}G_{v}^{\circ}(\overline{\mathbb{Q}}_{p})-$conjugacy class $\mathrm{Sat}(\pi_{v})$, the Satake parameter of $\pi_{v}$.
@@ A.1. Integral models of PEL-type Shimura varieties
We keep the notations and assumptions of section 2. Let $G$ be the connected reductive group defined by the rational PEL datum and $\mathscr{L}$ a multichain of $\operatorname{\mathcal{O}}_{B}$-lattices (see Definition A.1). By [49], there are three integral models $\mathscr{A}^{\operatorname{naive}},\mathscr{A}$ and $\mathscr{A}^{\mathrm{spl}}$ of the PEL moduli space over $E$ (with respect to the multichain $\mathscr{L}$). The models $\mathscr{A}^{\operatorname{naive}},\mathscr{A}$ are defined over $\operatorname{\mathcal{O}}_{E}$ and $\mathscr{A}^{\mathrm{spl}}$ is defined over $\operatorname{\mathcal{O}}_{F}$. We first recall the definition of PEL datum with parahoric level structure following [55], see also the appendix of [59] or [23] section 2.
@@ A.1.1. Parahoric data at $p$
    A chain of $\operatorname{\mathcal{O}}_{B}$-lattices in $V$ is a set of totally ordered $\operatorname{\mathcal{O}}_{B}$-lattices $\mathscr{L}$ such that for every element $x\in B^{\times}$ which normalizes $\operatorname{\mathcal{O}}_{B}$, one has
@@
for the group scheme over $\mathbb{Z}_{p}$ defined by the stabilizer of the multichain $\mathscr{L}$ as in [55]*§6. By [21]*Corollary 4.8, for a reductive group of the form $G=\operatorname{Res}_{F|\operatorname{\mathbb{Q}}_{p}}G^{\prime}$, every parahoric group scheme of $G$ is of the form $\operatorname{Res}_{\operatorname{\mathcal{O}}_{F}|\operatorname{\mathbb{Z}}_{p}}\operatorname{\mathcal{G}}^{\prime}$ for a unique parahoric group scheme $\operatorname{\mathcal{G}}^{\prime}$ of $G^{\prime}$. If we assume that $G^{\prime}$ unramified, the stabilizer group scheme $\operatorname{\mathcal{G}}=\operatorname{\mathcal{G}}_{\mathscr{L}}$ associated to $\mathscr{L}$ is always connected. So $\operatorname{\mathcal{G}}$ is a parahoric group scheme of $G$ and (since the prime to $p$ level $K^{p}$ is fixed) we write $K=\mathcal{G}(\mathbb{Z}_{p})$.
@@ A.1.2. Integral models
    ${{A_{\Lambda}^{a}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\theta_{a,\Lambda}}$${{A_{a\Lambda}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho_{a\Lambda,\Lambda}}$${{A_{\Lambda}}}$
@@
  • $\overline{\eta}$ is a $\pi_{1}(S,\overline{s})$-invariant $K^{p}$-orbit of isomorphism $\eta:V\otimes\mathbb{A}_{f}^{p}\to T^{p}(A_{\overline{s}})$ which preserves the pairings up to a scalar in $(\mathbb{A}_{f}^{p})^{\times}$. Here $T^{p}(A_{\overline{s}})$ is the prime to $p$ Tate module of $A_{\overline{s}}$ and for simplicity we assume that $S$ is connected.
@@
Suppose that $S$ is a scheme over $\operatorname{\mathcal{O}}_{F}$. A $\mathscr{L}$-set of polarized $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{S}$-modules is a triple $(\underline{\operatorname{\mathcal{H}}},\underline{\mathscr{F}},\underline{\iota})$, where:
@@
    $\underline{\operatorname{\mathcal{H}}}:\Lambda\mapsto\operatorname{\mathcal{H}}_{\Lambda}$ and $\underline{\mathscr{F}}:\Lambda\mapsto\mathscr{F}_{\Lambda}$ are functors from the category $\mathscr{L}$ to the category of $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{S}$-modules.
@@
    For each $\Lambda\in\mathscr{L}$, both $\mathscr{F}_{\Lambda}$ and $\operatorname{\mathcal{H}}_{\Lambda}/\mathscr{F}_{\Lambda}$ are finite locally free $\operatorname{\mathcal{O}}_{S}$-modules, such that $\operatorname{\mathcal{H}}_{\Lambda}/\mathscr{F}_{\Lambda}$ satisfies the determinant condition.
@@
Suppose that $S$ is a scheme over $\operatorname{\mathcal{O}}_{F}$, and that $(\underline{\mathcal{H}},\underline{\mathscr{F}},\underline{\iota})$ is a $\mathscr{L}$-set of polarized $\operatorname{\mathcal{O}}_{B}\otimes\mathcal{O}_{S}$-modules. A splitting structure for $(\underline{\mathcal{H}},\underline{\mathscr{F}},\underline{\iota})$ is a collection
$\underline{\operatorname{\mathscr{F}}_{\bullet}}=(\mathscr{F}_{i,j}^{l},\iota_{i,j}^{l})_{1\leq i\leq r,1\leq j\leq f_{i},0\leq l\leq e_{i}},$
where each $\mathscr{F}_{i,j}^{l}:\Lambda\to\mathscr{F}_{\Lambda,i,j}^{l}$ is a functor from the category $\mathscr{L}$ to the category of $\operatorname{\mathcal{O}}_{B}\otimes\mathcal{O}_{S}$-modules, and each $\iota_{i,j}^{l}:\mathscr{F}_{\Lambda,i,j}^{l}\to\operatorname{\mathcal{H}}_{\Lambda,i,j}$ is an injective morphism satisfying the following conditions (identify $\mathscr{F}_{i,j}^{l}$ with its image under $\iota$):
@@
    For each $\Lambda\in\mathscr{L}$, we require both $\mathscr{F}_{\Lambda,i,j}^{l}$ and $\operatorname{\mathcal{H}}_{\Lambda,i,j}/\mathscr{F}_{\Lambda,i,j}^{l}$ to be finite locally free $\mathcal{O}_{S}$-modules.
@@
    For each $\Lambda\in\mathscr{L}$ and tuples $(i,j,l)$, let $(\mathscr{F}_{\Lambda,i,j}^{l})^{\perp}$ denote the orthogonal complement of $\mathscr{F}_{\Lambda,i,j}^{l}$ in $\operatorname{\mathcal{H}}_{\Lambda^{\vee},i,j}$ with respect to the perfect pairing $\operatorname{\mathcal{H}}_{\Lambda,i,j}\times\operatorname{\mathcal{H}}_{\Lambda^{\vee},i,j}\to\operatorname{\mathcal{O}}_{S}$. Then
@@
Let $\mathscr{A}^{\mathrm{spl}}_{\mathscr{L}}$ be the moduli scheme over $\operatorname{\mathcal{O}}_{F}$ classifying the objects $(A_{\mathscr{L}},\overline{\lambda},\overline{\eta},\underline{\mathscr{F}_{\bullet}})$ for each scheme $S/\operatorname{\mathcal{O}}_{F}$, where $(A_{\mathscr{L}},\overline{\lambda},\overline{\eta})$ is an object in $\mathscr{A}^{\operatorname{naive}}_{\mathscr{L}}(S)$ and $\underline{\mathscr{F}_{\bullet}}$ is a splitting structure for the $\mathscr{L}$-set of polarized $\operatorname{\mathcal{O}}_{B}\otimes\operatorname{\mathcal{O}}_{S}$-modules associated with $(A_{\mathscr{L}},\overline{\lambda},\overline{\eta})$. This is the splitting model associated to the PEL datum. With the notation of [37], we have
@@ A.2. Local models
The local structure of integral models of Shimura varieties is controlled by the associated local models. There is a naive local model $\mathbb{M}^{\operatorname{naive}}$ associated to the moduli scheme $\mathscr{A}^{\operatorname{naive}}$. Recall that $\operatorname{\mathbb{M}}^{\operatorname{naive}}:=\operatorname{\mathbb{M}}^{\operatorname{naive}}(\mathcal{G},\mu)$ is the moduli scheme over $\operatorname{\mathcal{O}}_{E}$ given by the following definition.
@@
[…]
@@
We require the following conditions to be satisfied:
@@
    The morphisms $\varphi_{\Lambda}$ are surjective.
@@
as the moduli scheme over $\operatorname{\mathcal{O}}_{F}$ classifying the splitting structures over $\mathbb{M}^{\operatorname{naive}}\otimes_{\operatorname{\mathcal{O}}_{E}}\operatorname{\mathcal{O}}_{F}$. The splitting model $\mathbb{M}^{\mathrm{spl}}$ admits good properties. Recall we assume that there is no type (AR) local factors.
@@
Such a multichain determines a parahoric subgroup of $G^{\prime}_{\tau}$. This gives a local model (over $\operatorname{\mathcal{O}}_{F}$)
@@
By our assumption, $G^{\prime}$ is unramified over $F$, so $\operatorname{\mathbb{M}}^{\mathrm{loc}}$ is the same as the naive local model. This means that for each $\operatorname{\mathcal{O}}_{F}$-scheme $S$, $\operatorname{\mathbb{M}}^{\mathrm{loc}}(S)$ classifies the set of multichains $\{\operatorname{\mathscr{F}}_{\Lambda,\tau}\}_{\Lambda\in\operatorname{\mathscr{L}}}\subset\operatorname{\mathscr{L}}_{\tau}$ which is compatible with transition maps, and each $\operatorname{\mathscr{F}}_{\Lambda,\tau}$ is Zariski locally on $S$ an $\operatorname{\mathcal{O}}_{S}$-direct summand of $\Lambda_{\tau}$ of rank $d_{\tau}$.
@@
Such a trivialization exists Zariski locally on $S$ by [49]*Propositions 5.2, 9.2 (this reference only covers the case of type (C) and (AL); the case of type (AU) can be proved by the same argument). The map $\pi_{1}$ is the natural forgetful morphism, and $\pi_{2}$ sends $(\operatorname{\mathscr{F}}_{\Lambda,j}^{l},{\varphi_{\Lambda,j}^{l}})$ to $\varphi_{\Lambda,j}^{l}(\operatorname{\mathscr{F}}_{\Lambda,j}^{l}/\operatorname{\mathscr{F}}_{\Lambda,j}^{l-1})\in\operatorname{\mathbb{M}}^{\mathrm{loc}}_{\sigma_{j}^{l}}$.
[…]
@@
[…]
@@
We only need to show that $\operatorname{\mathbb{M}}$ coincides with $\operatorname{\mathbb{M}}^{\rm loc}$, as the other statements are contained in [49]. The image $\operatorname{\mathbb{M}}^{\rm loc}$ is flat by the flatness of $\operatorname{\mathbb{M}}^{\mathrm{spl}}$. As the generic fiber of $\operatorname{\mathbb{M}}$ agrees with $\operatorname{\mathbb{M}}^{\operatorname{naive}}$, $\operatorname{\mathbb{M}}$ is the flat closure of the generic fiber of $\operatorname{\mathbb{M}}^{\operatorname{naive}}$, which agrees with $\operatorname{\mathbb{M}}^{\rm\mathrm{loc}}$ by definition. ∎
@@
For each $i,j,l$, let $\operatorname{\mathcal{G}}_{i,j}^{l}$ be the group scheme over $\operatorname{\mathcal{O}}_{F}$ defined by the automorphism of the multichain $(\Lambda_{i,j}^{l})_{\Lambda\in\operatorname{\mathscr{L}}}$. Then it admits a decomposition as $\operatorname{\mathcal{G}}^{\mathrm{spl}}=\prod_{i,j,l}\mathcal{G}^{l}_{i,j}$. Now we define
@@
where $\pi^{\prime\prime}$ is a $\operatorname{\mathcal{G}}_{\operatorname{\mathcal{O}}_{F}}\times\prod_{i,j,l\geq 2}\operatorname{\mathcal{G}}_{i,j}^{l}$-torsor, and we can check that the same action of $\pi^{\prime\prime}$ makes $q^{\prime\prime\prime}$ a $\operatorname{\mathcal{G}}_{\operatorname{\mathcal{O}}_{F}}\times\prod_{i,j,l\geq 2}\operatorname{\mathcal{G}}_{i,j}^{l}$-equivariant smooth morphism. Note that we always have $\Lambda_{i}^{1}\subset\Lambda$, so the restriction of the $\operatorname{\mathcal{G}}$-action on $\Lambda$ induces a natural morphism